OpenStudy (anonymous):

Eric drove 375 miles using 14 gallons of gas.At this rate, how many gallons of gas would he need to drive 300 miles?

OpenStudy (anonymous):

375/14 = 300/x (375 miles to 14 gal. = 300 miles to x gallons) now cross multiply... (375)(x) = (14)(300) 375x = 4200 x = 4200/375 x = 11.2 He would use 11.2 gallons to drive 300 miles.